A Financial Analysis has been studying the Mean Corporation and has decided that the probability that the company's stock price will go up on next monday is 0.4, the probability that the stock price will go up on tuesday is 0.6 and the probability that the stock price will go up on BOTH days is 0.2. Assume that on each day the stock price will either increase or decrease.
A) Find the probability that the stock price will go up on Monday but not on Tuesday.
B) Find the probability that the stock price will go up on tuesday but not on monday.
THANK YOU

Answer by venugopalramana(3286)

A Financial Analysis has been studying the Mean Corporation and has decided that the probability that the company's stock price will go up on next monday is 0.4, the probability that the stock price will go up on tuesday is 0.6 and the probability that the stock price will go up on BOTH days is 0.2. Assume that on each day the stock price will either increase or decrease.
A) Find the probability that the stock price will go up on Monday but not on Tuesday.
P GOING UP ON MONDAY = 0.4
P NOT GOING UP ON TUESDAY =1-0.6=0.4
P GOING UP ON MONDAY AND NOT GOING UP ON TUES DAY = 0.4*0.4=0.16
B) Find the probability that the stock price will go up on tuesday but not on monday.
P NOT GOING UP ON MONDAY =1-0.4=0.6
P GOING UP ON TUES DAY =0.6
COMBINED P = 0.6*0.6=0.36
THANK YOU